# The Boolean spectrum of a Grothendieck category

### v1

@@ Abstract.
A notion of support for objects in any Grothendieck category is introduced. This is based on the spectral category of a Grothendieck category and uses its Boolean lattice of localising subcategories. The support provides a classification of all subcategories that are closed under arbitrary coproducts, subobjects, and essential extensions. There is also a notion of exact support which classifies certain thick subcategories. As an application we extend for any ring Crawley-Boevey’s correspondence between definable subcategories of modules and closed subsets of the Ziegler spectrum.
@@
Definable subcategory, Grothendieck category, spectral category, support, Ziegler spectrum
@@ 1. Introduction
The spectrum $\operatorname{Sp}\mathcal{A}$ of a Grothendieck category $\mathcal{A}$ is the set of isomorphism classes of indecomposable injective objects. This definition is due to Gabriel, and he uses this spectrum when he reconstructs a noetherian scheme from its category of quasi-coherent sheaves [10]. A finer invariant is introduced in work of Gabriel and Oberst [11]. The spectral category $\operatorname{Spec}\mathcal{A}$ is again a Grothendieck category with the additional propery that every short exact sequence is split exact. There is a canonical functor $\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ which identifies $\operatorname{Sp}\mathcal{A}$ with the isomorphism classes of simple objects in $\operatorname{Spec}\mathcal{A}$.
@@
(1.1) $\mathcal{A}\supseteq\mathcal{C}\longmapsto\operatorname{supp}(\mathcal{C})=\bigcup_{X\in\mathcal{C}}\operatorname{supp}(X)$
@@
A useful property of this support is that $\operatorname{supp}(X)=\varnothing$ implies $X=0$. This one cannot expect when using $\operatorname{Sp}\mathcal{A}$ as target for a support function, because $\operatorname{Sp}\mathcal{A}$ may be empty. Another option for defining support would be to use as target the frame $\mathbf{L}(\mathcal{A})$ of localising subcategories of $\mathcal{A}$. However, its set of points may be empty as well [34]. The following result shows that $\operatorname{Spc}\mathcal{A}$ is an invariant which encompasses both $\operatorname{Sp}\mathcal{A}$ and $\mathbf{L}(\mathcal{A})$.
@@
The work of Gabriel and his use of the spectrum of indecomposable injectives has been a great inspiration for noncommutative algebraic geometry; see for example [2, 19, 32] where several possible definitions of a spectrum of an abelian category are discussed. Here we follow a somewhat different route. For not necessarily commutative rings there is a spectrum which receives much attention and provides another motivation for this work.
@@
Let us give a quick outline of this paper. In §2 we recall the basic facts about spectral categories and introduce the Boolean spectrum of a Grothendieck category. In §3 we set up the support for Grothendieck categories and establish the classification of essentially closed subcategories. The following §4 provides the derived analogue of the results from §3 which leads to the exact support. The final §5 discusses the applications for module categories, but in the more general setting of exactly definable categories.
@@ 2. Spectral Grothendieck categories
We write $\operatorname{Spec}\mathcal{A}$ for the spectral category of $\mathcal{A}$ which is obtained from $\mathcal{A}$ by formally inverting all essential monomorphisms; see [11] or [33, V.7]. The category $\operatorname{Spec}\mathcal{A}$ is again a Grothendieck category in which all exact sequences are split exact. A Grothendieck category with this property is called spectral. The canonical functor
@@
maps any injective enevelope $X\to E(X)$ in $\mathcal{A}$ to an isomorphism and it identifies the indecomposable injective objects in $\mathcal{A}$ with the simple objects in $\operatorname{Spec}\mathcal{A}$. These properties are clear from the definition. Also, $P$ is left exact and preserves all coproducts since the class of essential monomorphisms is closed under all coproducts. More precisely, for any directed union $X=\sum_{\alpha}X_{\alpha}$ in $\mathcal{A}$ we have $P(X)=\sum_{\alpha}P(X_{\alpha})$. We write $\operatorname{Sp}\mathcal{A}$ for the set of isomorphism classes of indecomposable injective objects in $\mathcal{A}$, and $P$ induces a bijection $\operatorname{Sp}\mathcal{A}\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Sp}(\operatorname{Spec}\mathcal{A})$.
@@
where $\operatorname{Rad}\mathcal{C}$ denotes the Jacobson radical of an additive category $\mathcal{C}$; see [24, Proposition 2.5.9]. Recall that $\operatorname{Rad}\mathcal{C}$ is an ideal of morphisms that is given by subgroups
@@
where $E(X)$ denotes an injective envelope of $X$; see also [11, Satz 1.5].
The following lemma shows that the spectral category controls the decomposition of injective objects.
@@
Spectral categories and their structure are closely related to (modules over) self-injective von Neumann regular rings, because the spectral categories are precisely (up to equivalence) the categories of nonsingular injective modules over such rings. This is explained in [11, 30], and for the structure of such rings we refer to [13]. [^1] Given a spectral category and a generator $G$, the endomorphism ring of $G$ is self-injective von Neumann regular, and the equivalence is induced by $\operatorname{Hom}(G,-)$. There is an elaborated decomposition theory for operator algebras which uses certain types and goes back to Murray, von Neumann, and Kaplansky [20, 26]. The analogous decomposition theory for spectral categories is developed in [14, 30]. For some further structure of spectral categories, see [8].
@@
A complete lattice is a frame if finite meets distribute over arbitrary joins, and it is a coframe if finite joins distribute over arbitrary meets. We record the following well known fact; see for instance [12] for the case of a module category, and we refer to [34] for a proof that uses an explicit construction of the join in the lattice of localising subcategories.
@@
This yields a decomposition $\mathcal{A}=\mathcal{A}_{\mathrm{d}}\times\mathcal{A}_{\mathrm{c}}$ of spectral Grothendieck categories such that $\mathcal{A}_{\mathrm{d}}$ is discrete and $(\mathcal{A}_{\mathrm{c}})_{\mathrm{d}}=0=(\mathcal{A}_{\mathrm{d}})_{\mathrm{c}}$; see [11, §3].
@@
Let $\mathcal{A}$ be a locally noetherian Grothendieck category. Then its spectral category $\operatorname{Spec}\mathcal{A}$ is discrete [10, IV.2]. Thus in this case $\mathbf{L}(\operatorname{Spec}\mathcal{A})$ identifies with the lattice of subsets of $\operatorname{Sp}\mathcal{A}$.
@@
Localising subcategories of $\mathcal{A}$ are in bijective correspondence to hereditary torsion pairs, by taking a subcategory $\mathcal{C}$ to the pair $(\mathcal{C},\mathcal{C}^{\perp})$; see [33, VI.3]. Torsion pairs for spectral categories admit the following elementary description.
@@
Remark 2.10.
It would be interesting to know which Boolean lattices arise from spectral categories. More specifically, given a Boolean lattice $L$, is there a procedure to construct a self-injective von Neumann regular ring $\Lambda$ such that for the category $\mathcal{A}$ of singular injective $\Lambda$-modules we have $\mathbf{L}(\mathcal{A})\cong L$? There is an elaborated structure theory for self-injective von Neumann regular rings, which seems to neglect procedures for actual constructions.
@@
We briefly discuss the functoriality of the assignment $\mathcal{A}\mapsto\operatorname{Spec}\mathcal{A}$. An exact and coproduct preserving functor $\mathcal{A}\to\mathcal{B}$ between Grothendieck categories admits a right adjoint which restricts to an additive functor $\operatorname{Inj}\mathcal{B}\to\operatorname{Inj}\mathcal{A}$. Using (2.1) one obtains an exact functor $\operatorname{Spec}\mathcal{B}\to\operatorname{Spec}\mathcal{A}$.
@@
For a pair of objects $X,Y$ in $\mathcal{A}$ we have $\langle X\rangle\cap\langle X\rangle=\langle X\otimes_{G}Y\rangle$.
@@
We have $t_{\langle X\rangle}=t_{\langle X\rangle}(G)\otimes_{G}-$. Thus for an object $M\in\langle X\rangle\cap\langle X\rangle$ we have
@@ 3. Essentially closed subcategories
We begin with the main definition of this section; the terminology is justified by Theorem 3.3, because we look at the operations that are preserved by the functor $\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ inverting all essential monomorphisms.
@@
(1) $\Rightarrow$ (2): We set $\mathcal{C}:={{}^{\perp}\mathcal{D}}$ and this is localising since $\mathcal{D}$ is closed under injective envelopes, so $\mathcal{C}={{}^{\perp}(\mathcal{D}\cap\operatorname{Inj}\mathcal{A})}$. It remains to note that $\mathcal{C}^{\perp}=\mathcal{D}$ where one uses that $\mathcal{D}$ is closed under arbitrary products; see [33, Proposition VI.2.2].
@@
This is an immediate consequence of Proposition 3.6. ∎
@@
It is easily checked that the essential extensions of objects in $\mathcal{C}$ form an essentially closed subcategory. Now let $(\mathcal{C},\mathcal{D})$ be a hereditary torsion pair for $\mathcal{A}$. From Lemmas 2.7 and 3.16 it follows that $(P(\bar{\mathcal{C}}),P(\mathcal{D}))$ is a hereditary torsion pair for $\operatorname{Spec}\mathcal{A}$. ∎

### v2

@@ Abstract.
A notion of support for objects in any Grothendieck category is introduced. This is based on the spectral category of a Grothendieck category and uses its Boolean lattice of localising subcategories. The support provides a classification of all subcategories that are closed under arbitrary coproducts, subobjects, and essential extensions. There is also a notion of exact support which classifies certain thick subcategories. As an application, the coproduct decompositions of objects are described in terms of Boolean lattices. Also, for any ring Crawley-Boevey’s correspondence between definable subcategories of modules and closed subsets of the Ziegler spectrum is extended.
@@
Boolean lattice, definable subcategory, Grothendieck category, spectral category, support, Ziegler spectrum
@@ 1. Introduction
The spectrum $\operatorname{Sp}\mathcal{A}$ of a Grothendieck category $\mathcal{A}$ is the set of isomorphism classes of indecomposable injective objects. This definition is due to Gabriel, and he uses this spectrum when he reconstructs a noetherian scheme from its category of quasi-coherent sheaves [12]. A finer invariant is introduced in work of Gabriel and Oberst [13]. The spectral category $\operatorname{Spec}\mathcal{A}$ is again a Grothendieck category with the additional property that every short exact sequence is split exact. There is a canonical functor $\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ which identifies $\operatorname{Sp}\mathcal{A}$ with the isomorphism classes of simple objects in $\operatorname{Spec}\mathcal{A}$.
@@
(1.1) $\mathcal{C}\longmapsto\operatorname{supp}(\mathcal{C})=\bigvee_{X\in\mathcal{C}}\operatorname{supp}(X)$
@@
The local structure of an object can be described in terms of its support. More precisely, we introduce for each injective object $X$ a complete Boolean lattice $\mathbf{D}(X)$ which is given by the clopen subsets of $\operatorname{supp}(X)$; its elements parameterise the direct summands of $X$ up to an appropriate equivalence relation and there is a canonical isomorphism
$\mathbf{D}(X)\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{B}(\operatorname{End}(X)/J(\operatorname{End}(X))),$
where $\mathbf{B}(\Lambda)$ denotes the lattice of central idempotents of a ring $\Lambda$ (Theorem 4.5). The ring $\operatorname{End}(X)/J(\operatorname{End}(X))$ is von Neumann regular; so an elaborated theory of decompositions and types applies [23, 29]. It is interesting to note that any complete Boolean lattice can be realised as $\mathbf{D}(X)$ for some injective object $X$ and some appropriate Grothendieck category.
The support of an object has the useful property that $\operatorname{supp}(X)=\varnothing$ implies $X=0$. This one cannot expect when using $\operatorname{Sp}\mathcal{A}$ as target for a support function, because $\operatorname{Sp}\mathcal{A}$ may be empty. Another option for defining support would be to use as target the frame $\mathbf{L}(\mathcal{A})$ of localising subcategories of $\mathcal{A}$. However, its set of points may be empty as well [37]. The following result shows that $\operatorname{Spc}\mathcal{A}$ is an invariant which encompasses both $\operatorname{Sp}\mathcal{A}$ and $\mathbf{L}(\mathcal{A})$.
@@
The work of Gabriel and his use of the spectrum of indecomposable injectives has been a great inspiration for non-commutative algebraic geometry; see for example [3, 22, 35] where several possible definitions of a spectrum of an abelian category are discussed. Here we follow a somewhat different route. For not necessarily commutative rings there is a spectrum which receives much attention and provides another motivation for this work.
@@
Let us give a quick outline of this paper. In §2 we recall the basic facts about spectral categories and introduce the Boolean spectrum of a Grothendieck category. In §3 we set up the support for Grothendieck categories and establish the classification of essentially closed subcategories. A first application is given in §4, where coproduct decompositions of objects are studied in terms of Boolean lattices. The following §5 provides the derived analogue of the results from §3, which leads to the exact support. The final §6 discusses the applications for module categories, but in the more general setting of exactly definable categories.
@@ 2. Spectral Grothendieck categories
We write $\operatorname{Spec}\mathcal{A}$ for the spectral category of $\mathcal{A}$ which is obtained from $\mathcal{A}$ by formally inverting all essential monomorphisms; see [13] or [36, V.7]. The category $\operatorname{Spec}\mathcal{A}$ is again a Grothendieck category in which all exact sequences are split exact. A Grothendieck category with this property is called spectral. The canonical functor
@@
maps any injective envelope $X\to E(X)$ in $\mathcal{A}$ to an isomorphism and it identifies the indecomposable injective objects in $\mathcal{A}$ with the simple objects in $\operatorname{Spec}\mathcal{A}$. These properties are clear from the definition. Also, $P$ is left exact and preserves all coproducts since the class of essential monomorphisms is closed under all coproducts. More precisely, for any directed union $X=\sum_{\alpha}X_{\alpha}$ in $\mathcal{A}$ we have $P(X)=\sum_{\alpha}P(X_{\alpha})$. We write $\operatorname{Sp}\mathcal{A}$ for the set of isomorphism classes of indecomposable injective objects in $\mathcal{A}$, and $P$ induces a bijection $\operatorname{Sp}\mathcal{A}\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Sp}(\operatorname{Spec}\mathcal{A})$.
@@
where $\operatorname{Rad}\mathcal{C}$ denotes the Jacobson radical of an additive category $\mathcal{C}$; see [27, Proposition 2.5.9]. Recall that $\operatorname{Rad}\mathcal{C}$ is an ideal of morphisms that is given by subgroups
@@
where $E(X)$ denotes an injective envelope of $X$; see also [13, Satz 1.5].
The following lemma shows that the spectral category controls the decompositions of injective objects.
@@
Spectral categories and their structure are closely related to (modules over) self-injective von Neumann regular rings, because the spectral categories are precisely (up to equivalence) the categories of non-singular injective modules over such rings. This is explained in [13, 33], and for the structure of such rings we refer to [15]. [^1] Given a spectral category and a generator $G$, the endomorphism ring of $G$ is self-injective von Neumann regular, and the equivalence is induced by $\operatorname{Hom}(G,-)$. There is an elaborated decomposition theory for operator algebras which uses certain types and goes back to Murray, von Neumann, and Kaplansky [23, 29]. The analogous decomposition theory for spectral categories is developed in [16, 33]. For some further structure of spectral categories, see [10].
@@
A complete lattice is a frame if finite meets distribute over arbitrary joins, and it is a coframe if finite joins distribute over arbitrary meets. It is well known that the localising subcategories of a Grothendieck form a frame; see for instance [2] (in a more general categorical context) or [14] (for module categories). We refer to [37] for a proof that uses an explicit construction of the join in the lattice of localising subcategories.
@@
This yields a decomposition $\mathcal{A}=\mathcal{A}_{\mathrm{d}}\times\mathcal{A}_{\mathrm{c}}$ of spectral Grothendieck categories such that $\mathcal{A}_{\mathrm{d}}$ is discrete and $(\mathcal{A}_{\mathrm{c}})_{\mathrm{d}}=0=(\mathcal{A}_{\mathrm{d}})_{\mathrm{c}}$; see [13, §3].
@@
Let $\mathcal{A}$ be a locally noetherian Grothendieck category. Then its spectral category $\operatorname{Spec}\mathcal{A}$ is discrete [12, IV.2]. Thus in this case $\mathbf{L}(\operatorname{Spec}\mathcal{A})$ identifies with the lattice of subsets of $\operatorname{Sp}\mathcal{A}$.
@@
Localising subcategories of $\mathcal{A}$ are in bijective correspondence to hereditary torsion pairs, by taking a subcategory $\mathcal{C}$ to the pair $(\mathcal{C},\mathcal{C}^{\perp})$; see [36, VI.3]. Torsion pairs for spectral categories admit the following elementary description.
@@
Central idempotents
For an additive category $\mathcal{A}$ let $Z(\mathcal{A})$ denote its centre which is the ring of natural transformations $\operatorname{id}_{\mathcal{A}}\to\operatorname{id}_{\mathcal{A}}$. For a ring $\Lambda$ let $\mathbf{B}(\Lambda)$ denote the lattice of central idempotents, with operations given by
$x\vee y:=x+y-x\cdot y\qquad\text{and}\qquad x\wedge y:=x\cdot y.$
Let $\mathcal{A}$ be a spectral Grothendieck category. Any torsion pair $(\mathcal{C},\mathcal{D})$ gives rise to a functorial decomposition $X=t_{\mathcal{C}}(X)\oplus t_{\mathcal{D}}(X)$ for each object $X\in\mathcal{A}$, which amounts to an idempotent $e_{\mathcal{C}}\colon\operatorname{id}_{\mathcal{A}}\to\operatorname{id}_{\mathcal{A}}$ such that the image of $(e_{\mathcal{C}})_{X}$ equals $t_{\mathcal{C}}(X)$. Conversely, any idempotent $e\in Z(\mathcal{A})$ yields a torsion pair $(\mathcal{C},\mathcal{D})$ if we set $\mathcal{C}=\{X\in\mathcal{A}\mid\operatorname{Im}e_{X}=X\}$ and $\mathcal{D}=\{X\in\mathcal{A}\mid\operatorname{Ker}e_{X}=X\}$. Clearly, these operations are mutually inverse to each other, and this yields another description of $\mathbf{L}(\mathcal{A})$.
Lemma 2.10.
Let $\mathcal{A}$ be a spectral Grothendieck category. Then the assignment $\mathcal{C}\mapsto e_{\mathcal{C}}$ induces a lattice isomorphism
$\mathbf{L}(\mathcal{A})\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{B}(Z(\mathcal{A})).$
Moreover, when $G\in\mathcal{A}$ is a generator with $\Lambda=\operatorname{End}(G)$, then
$x\longmapsto\{X\in\mathcal{A}\mid\operatorname{Hom}(G,X)x=\operatorname{Hom}(G,X)\}$
induces a lattice isomorphism
$\mathbf{B}(\Lambda)\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{L}(\mathcal{A}).$
Proof.
The first assertion is clear from the above discussion. Let $\mathcal{A}^{\prime}$ denote the category of non-singular injective right modules over $\Lambda$. The functor $\operatorname{Hom}(G,-)$ induces an equivalence $\mathcal{A}\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathcal{A}^{\prime}$ and an isomorphism $Z(\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}Z(\mathcal{A}^{\prime})$. On the other hand, multiplication with a central element induces an isomorphism $Z(\Lambda)\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}Z(\mathcal{A}^{\prime})$. Thus the second assertion follows from the first. ∎
I am grateful to Ken Goodearl for pointing out the following converse to Proposition 2.8.
Proposition 2.11 (Goodearl).
Every complete Boolean lattice is isomorphic to $\mathbf{L}(\mathcal{A})$ for some spectral Grothendieck category $\mathcal{A}$.
Proof.
We begin with the following observation. Let $\Lambda$ be a right self-injective von Neumann regular ring and let $\mathcal{A}$ denote the corresponding spectral category consisting of the non-singular injective $\Lambda$-modules. Then it follows from Lemma 2.10 that the assignment $x\mapsto\{M\in\mathcal{A}\mid Mx=M\}$ induces a lattice isomorphism $\mathbf{B}(\Lambda)\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathbf{L}(\mathcal{A})$.
Now let $L$ be a complete Boolean lattice and view this as a Boolean ring. Then $L$ is von Neumann regular with $\mathbf{B}(L)=L$, and the ring is self-injective because $L$ is complete. Thus for $\mathcal{A}$ the category of non-singular injective $L$-modules, we have that $L\cong\mathbf{L}(\mathcal{A})$. ∎
The proof shows that the realisation of a complete Boolean lattice is canonical. For a ring $\Lambda$ let $\operatorname{Inj}_{\mathrm{ns}}\Lambda$ denote the category of non-singular injective right modules over $\Lambda$. Then for any complete Boolean lattice $L$ there is a canonical isomorphism
$L\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{L}(\operatorname{Inj}_{\mathrm{ns}}L).$
The above result suggests a procedure for constructing explicit examples of objects with specific decomposition properties. For instance, we obtain examples of superdecomposable objects. Recall that an object in an abelian category is superdecomposable if there are no indecomposable direct summands.
Example 2.12 (Goodearl).
Let $\Lambda$ be an infinite direct product of copies of $\mathbb{Z}/2$ and let $\bar{\Lambda}$ denote the maximal quotient ring of $\Lambda/\operatorname{soc}(\Lambda)$. Then $\bar{\Lambda}$ is injective as a $\bar{\Lambda}$-module and superdecomposable by construction, since $\operatorname{soc}(\bar{\Lambda})=0$. Moreover, $\mathbf{B}(\bar{\Lambda})=\bar{\Lambda}$ for its Boolean lattice of direct summands.
The ring $\bar{\Lambda}$ has characteristic $2$ and is of type I, but further examples of different characteristic and different type can be constructed via methods from Propositions 5-3.11 and 5-3.12 in [17].
For general Grothendieck categories the decompositions of objects will be discussed in more detail in §4.
@@
We briefly discuss the functoriality of the assignment $\mathcal{A}\mapsto\operatorname{Spec}\mathcal{A}$. An exact and coproduct preserving functor $\mathcal{A}\to\mathcal{B}$ between Grothendieck categories admits a right adjoint which restricts to an additive functor $\operatorname{Inj}\mathcal{B}\to\operatorname{Inj}\mathcal{A}$. Using (2.1) one obtains an exact functor $\operatorname{Spec}\mathcal{B}\to\operatorname{Spec}\mathcal{A}$, at least when $\mathcal{A}\to\mathcal{B}$ is the inclusion of a localising subcategory or a quotient functor.
@@
For a pair of objects $X,Y$ in $\mathcal{A}$ we have $\langle X\rangle\cap\langle Y\rangle=\langle X\otimes_{G}Y\rangle$.
@@
We have $t_{\langle X\rangle}=t_{\langle X\rangle}(G)\otimes_{G}-$. Thus for an object $M\in\langle X\rangle\cap\langle Y\rangle$ we have
@@ 3. Essentially closed subcategories
We begin with the main definition of this section; the terminology is justified by Theorem 3.3, because we look at the operations that are preserved by the functor $P\colon\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ inverting all essential monomorphisms.
@@
(1) $\Rightarrow$ (2): We set $\mathcal{C}:={{}^{\perp}\mathcal{D}}$ and this is localising since $\mathcal{D}$ is closed under injective envelopes, so $\mathcal{C}={{}^{\perp}(\mathcal{D}\cap\operatorname{Inj}\mathcal{A})}$. It remains to note that $\mathcal{C}^{\perp}=\mathcal{D}$ where one uses that $\mathcal{D}$ is closed under arbitrary products; see [36, Proposition VI.2.2].
@@
Consider $\mathcal{C}=\mathcal{C}_{1}\vee\mathcal{C}_{2}$. If $X_{1}\in\mathcal{C}_{1}$ and $X_{2}\in\mathcal{C}_{2}$, then $X_{1}\oplus X_{2}\in\mathcal{C}_{1}\vee\mathcal{C}_{2}$. Now assume $X\in\mathcal{C}_{1}\vee\mathcal{C}_{2}$. Set $Y=P(X)$ and $Y_{1}=t_{P(\mathcal{C}_{1})}(Y)$. Then $Y_{2}=Y/Y_{1}$ is in $P(\mathcal{C}_{2})$, and therefore $Y=Y_{1}\oplus Y_{2}$ with $Y_{i}\in P(\mathcal{C}_{i})$. Using Lemma 2.1 we find $X_{i}\in\mathcal{C}_{i}$ such that $X=X_{1}\oplus X_{2}$. See also Proposition 4.1. ∎
@@
It is easily checked that the essential extensions of objects in $\mathcal{C}$ form an essentially closed subcategory. Now let $(\mathcal{C},\mathcal{D})$ be a hereditary torsion pair for $\mathcal{A}$. From Lemmas 2.7 and 3.14 it follows that $(P(\bar{\mathcal{C}}),P(\mathcal{D}))$ is a hereditary torsion pair for $\operatorname{Spec}\mathcal{A}$. ∎
@@
4. Coproduct decompositions
Decompositions of objects in a Grothendieck have been studied in great detail in the past. For instance, the theorems of Azumaya, Krull, Remak, and Schmidt treat the decompositions into indecomposable objects with local endomorphism rings. [^3] Splittings or decompositions have been considered since ancient times. The Quran reports about a miraculous splitting of the moon [Quran 54:1-2], and the Bible about the splitting of the Red Sea [Exodus 14:21-22]. For arbitrary objects the decomposition theory is far more complex. In this section we discuss a Boolean lattice which controls the decompositions, at least for injective objects. There is probably no efficient way to describe all decompositions. So we focus on decompositions up to a certain equivalence relation, which reflects the decompositions of essentially closed subcategories.
We fix a Grothendieck category $\mathcal{A}$ and recall that $P\colon\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ denotes the functor that inverts all essential monomorphisms.
Coproduct decompositions via essentially closed subcategories
Proposition 4.1.
  1. (1)
  2. (2)
Proof.
Remark 4.2.
Remark 4.3.
In part (2) of the above proposition, the assumption on $X$ to be injective cannot be removed. Consider for instance an indecomposable object $X$ of length $3$ with $\operatorname{soc}(X)=S_{1}\oplus S_{2}$ such that $S_{1}$ and $S_{2}$ are non-isomorphic simple objects. Then $\langle X\rangle=\langle S_{1}\rangle\vee\langle S_{2}\rangle$ and $\langle S_{1}\rangle\wedge\langle S_{2}\rangle=0$, but there is no decomposition $X=X_{1}\oplus X_{2}$ with $\langle X_{i}\rangle=\langle S_{i}\rangle$.
The lattice of decompositions
We introduce an equivalence relation on the direct summands of an injective object in order to provide an efficient description of all decompositions.
Definition 4.4.
Let $X$ be an injective object in $\mathcal{A}$. Given a pair of direct summands $U\subseteq X$ and $V\subseteq X$, we say they are essentially equivalent if $\langle U\rangle=\langle V\rangle$. Let $\mathbf{D}(X)$ denote the set of essential equivalence classes of direct summands of $X$, with partial order given by $U\leq V$ if $\langle U\rangle\subseteq\langle V\rangle$.
The following result establishes a connection to the decomposition theory for von Neumann regular rings since $\operatorname{End}(X)/J(\operatorname{End}(X))$ has this property. In particular, the elaborated theory of types applies [23, 29].
Theorem 4.5.
Given an injective object $X$ in $\mathcal{A}$, the partially ordered set $\mathbf{D}(X)$ of direct summands is a complete Boolean lattice. Moreover, there is a canonical lattice isomorphism
$\mathbf{D}(X)\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{B}(\operatorname{End}(X)/J(\operatorname{End}(X))).$
Proof.
We claim that the assignment $U\mapsto\langle P(U)\rangle=P(\langle U\rangle)$ induces a lattice isomorphism
(4.1) $\mathbf{D}(X)\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\{\mathcal{U}\in\mathbf{S}(\mathcal{A})\mid\mathcal{U}\subseteq\operatorname{supp}(X)\}.$
The inverse map sends a localising subcategory $\mathcal{U}\subseteq\operatorname{Spec}\mathcal{A}$ to the equivalence class of a direct summand $U\subseteq X$ such that $P(U)=t_{\mathcal{U}}(P(X)$; for the existence of $U$ see Lemma 2.1. Using Theorem 3.3 it is easily checked that these assignments are mutually inverse to each other and preserving the partial order. Thus $\mathbf{D}(X)$ is a Boolean lattice, since $\mathbf{S}(\mathcal{A})$ has this property by Proposition 2.8. Moreover, the above argument shows that $P\colon\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ induces a lattice isomorphism
(4.2) $\mathbf{D}(X)\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{D}(P(X)).$
For the isomorphism with the lattice of central idempotents we consider the Grothendieck category
$\mathcal{A}_{X}:=\langle P(X)\rangle\subseteq\operatorname{Spec}\mathcal{A}$
which is generated by $P(X)$. The functor $P$ induces an isomorphism
$\operatorname{End}(X)/J(\operatorname{End}(X))\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\operatorname{End}(P(X))$
by (2.1). This yields the following sequence of isomorphisms
$\mathbf{D}(X)\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathbf{D}(P(X))\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathbf{L}(\mathcal{A}_{X})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathbf{B}(\operatorname{End}(X)/J(\operatorname{End}(X))),$
where the first is (4.2), the second is (4.1), and the last comes from Lemma 2.10. ∎
Remark 4.6.
The Boolean structure of $\mathbf{D}(X)$ implies that each equivalence class contains a summand $U\subseteq X$ that is maximal in the sense that $\langle U\rangle\cap\langle X/U\rangle=0$. An equivalent condition is that for any pair of monomorphisms $U\leftarrow V\to X/U$ we have $V=0$.